A tyre of mass 15kg is suspended on a rope of length 7.5m and a girl of mass 75kg sits in the tyre. If the girls speed at the lowest point as she swings is 5m/s and the rope extends by 20cm, find the radius of this rope(young modulus = 2.0 x 10^11 Pa)
1
Expert's answer
2021-02-02T09:32:23-0500

Answer

radius of rope is given by

"r=\sqrt{\frac{mgl}{\pi \Delta l Y}}\\=\sqrt{\frac{(15+75)\times10\times7.5}{3.14\times0.2\times 2.0 \times 10^{11} }}\\=231 \times10^{-6}m"